OpenStudy (anonymous):

A cone has a radius of 8 inches and a height of 12 inches. What is the exact volume of the cone?

OpenStudy (anonymous):

@iGreen @k12user

OpenStudy (unanimoose):

Use the formula to find the volume of a cone. \[V = r^2\pi\frac{h}{3}\]

OpenStudy (igreen):

\(V = \dfrac{1}{3} \pi r^2 h\)

OpenStudy (unanimoose):

Where 'r' is radius and 'h' is height.

OpenStudy (igreen):

\(V = \dfrac{1}{3} \pi (8)^2 (12)\) Multiply 1/3 * 8 * 8 * 12

OpenStudy (igreen):

That will give us our answer in terms of pi.

OpenStudy (unanimoose):

Although when I solve it myself it doesn't show as any of the options . . .

OpenStudy (anonymous):

uhh.....not sure how to.

OpenStudy (unanimoose):

Nvm I did it wrong.

OpenStudy (unanimoose):

I multiplied pi, heh.

OpenStudy (igreen):

Lol,

OpenStudy (igreen):

>.>

OpenStudy (unanimoose):

No the equation was right I was just applying pi when i realized you're not supposed to because the question just has options with pi int hem.

OpenStudy (igreen):

Yes, yes, that's why I deleted my comment :P

OpenStudy (anonymous):

does any one know?

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@BloomLocke367